a rectangle has a length 2m less than twice its width. When 3 m are added to the width, the resulting figures is a square with an area of 64 m^2. Find the dimensions of the original rectangle
Answer by lwsshak3(11628)

a rectangle has a length 2m less than twice its width. When 3 m are added to the width, the resulting figures is a square with an area of 64 m^2. Find the dimensions of the original rectangle
let x=width
2x-2=length
(x+3)^2=x^2+6x+9=64
x^2+6x-55=0
(x+11)(x-5)=0
x=5
2x-2=8
dimensions of the original rectangle:
width=5 m
length=8 m